# Wasserstein Flow Matching: Generative modeling over families of distributions

### v1

@@ 4.1.1 Toy Datasets
MERFISH (16D) $\mathbf{1.90}$ $1.98$ $2.06$
@@
Table 1: Average minimum $W^{2}_{2}$ distance between each generated Gaussian and the reference datasets. We benchmark our BW-FM algorithm with Riemannian updates against (1) BW-FM with Euclidean updates, and (2) Frobenius FM; see Figure 2. Despite identical training schemes, the BW-FM (R) vastly outperforms other approaches on both synthetic and real data.
@@ 4.1.2 Single-Cell Genomics
Spatial transcriptomics are a set of techniques which build on single-cell genomics and preserve physical information of cells’ location in tissues, while assaying their gene expression. Haviv et al., 2024b demonstrated that a cell’s microenvironment can be effectively characterized using the mean and covariance of the surrounding cells gene expression. This statistical representation efficiently captures the key features of cellular neighborhoods and transforms each spatial transcriptomics dataset into a measure within the Bures–Wasserstein space, highlighting the value of generative modeling in this context.
In the motor cortex, excitatory neurons form phenotypically distinct and highly specialized cortical layer (Zeng and Sanes,, 2017). From the 254 gene MERFISH atlas (Zhang et al.,, 2021), we compute the mean and covariance of the top $16$ principal components of gene expression from all cells within an $80$ micron radius around each microglia. Despite the dimensionality of this data, BW-FM synthesizes Gaussians which are highly congruent with the real data (Figure 3 and Figure S2).
@@
Another common instance of BW manifolds arising in single-cell genomics is through aggregating cells into common states. These clusters can be summarized by their mean gene expression and its covariance. On a scRNA-seq atlas elucidating human immune response to COVID (Stephenson et al.,, 2021), we combine cells into MetaCells (Persad et al.,, 2023), and quantify the gene expression mean and covariance for each. Here too we apply BW-FM conditioned on cell-state, which encompass the heterogeneity of immune profiles appearing as response to COVID infection. Despite the plurality of labels, BW-FM can still synthesize appropriate examples for each condition (see Figure S1).
4.2 Flow matching between families of point-clouds
When the two measures are point-clouds, we turn to entropic optimal transport to estimate the McCann interpolation (Cuturi,, 2013). When the dataset consists of source and target point-clouds of the same size, we use a GPU-efficient rounding scheme to approximate the OT map using the entropic OT coupling; see Appendix A. When the support size of the source and target point-clouds vary, it is worth mentioning that an OT map may not even exist. Nevertheless, we approximate these curves using the entropic transport map of Pooladian and Niles-Weed, (2021). Together, these two approaches method offer a computationally feasible solution while maintaining accuracy in transport map estimation.
MNIST ($4$) Letters ($A$) seqFISH XENIUM
CD $\downarrow$ EMD $\downarrow$ CD $\downarrow$ EMD $\downarrow$ CD $\downarrow$ EMD $\downarrow$ CD $\downarrow$ EMD $\downarrow$
Current methods NA NA NA NA NA NA NA NA
WFM (ours) 63.34 59.97 62.12 58.68 61.79 64.34 60.69 64.20
Table 2: 1-Nearest-Neighbour Accuracy for high-dimensional or variable size point-clouds. WFM employs a transformer backbone and relies on the efficient computation of the entropic transport map, allowing it to scale to arbitrary dimensions and learn flows between point-cloud of variable sizes, key features all previous point-cloud generation approaches lack.
We compare WFM to many other point-cloud generation algorithms. Following in their footsteps, we measure generation quality based on the 1-Nearest-Neighbour accuracy metric between generated and test-set point-clouds. On uniform, 3D datasets, WFM is competitive with current approaches (Table 3), but exemplifies itself with its unique ability to generate point-clouds with varying sizes and in high-dimensions (Table 2).
4.2.1 2D & 3D Point-Clouds
Figure 5: Generated point-clouds from MNIST and EMNIST datasets using WFM, where $n$ denotes the number of points in each cloud.
Derived from 3D CAD designs, ShapeNet & ModelNet (Wu et al.,, 2015; Chang et al.,, 2015) are touchstone point-cloud datasets in computational geometry. Trained individually on samples from the chair, car and plane classes of ShapeNet, WFM synthesized high quality point-clouds with diverse profiles and matches the performance of previous 3D generation algorithms; see Figure 4 and Table 3. Our framework’s versatility allows for seamless integration of label information during training, enabling the synthesis of point-clouds conditioned on specific classes. On the full 40-class ModelNet dataset, WFM learned condition dependent flows, allowing for the same initial point-cloud to generate a diverse cohort of shapes based on the desired label; see Figure 4. We stress that WFM is not restricted to only noisy source measures but can generate transformations between any two collections of point-clouds. To this end, we demonstrate that WFM can interpolate between two arbitrary elements in the dataset (e.g., between a lamp and a handbag) and complete the point-clouds based on partial profiles (e.g., generate the remaining parts of a plane); see Figure S3.
Another novel facet of WFM is its ability to perform generative modeling from inhomogeneous datasets, where the number of points varies between independent samples. This happens in the MNIST or Letters datasets, where data is generated by thresholding grayscale numerical values. In this setting, we use the entropic transport map to approximate the objective; see equation 14. WFM sets itself apart from other methods, which are restricted to uniform datasets, by leveraging the entropic OT map’s ability to compute feasible transformations between point-clouds of different sizes. Our experiments in Figure 5 demonstrate that WFM generates high-quality & diverse samples, despite large variability in the number of points per sample, which is itself a novel contribution.
@@ 4.2 Flow matching between families of point-clouds
WFM (ours) 73.45 71.72 58.98 57.77 56.53 57.95
Table 3: Using 1-Nearest-Neighbour Accuracy based on Earth Mover’s Distance (EMD) and Chamfer’s Distance (CD). Wasserstein Flow Matching is competitive with existing approaches (data from Wu et al., (2023)), while producing diverse samples, see Figure 4.
4.2.2 Spatial Transcriptomics (High Dimensional Point Clouds)
@@
In spatial transcriptomics, the niche of a cell is the point-cloud in high-dimensional gene-expression space of its immediate nearest neighbours. This approach is complementary to the BW representation of a niche (recall Section 4.1.2), and serves as a more high fidelity view suited for fine-grain interactions. Due to their high dimensionality, cellular microenvironments have remained beyond the reach of point-cloud-based generative models that depend on voxel-based neural networks. In contrast, WFM leverages transformers, which due to their permutation equivariance and indifference to dimensionality, have already emerged as possible architecture to model spatial transcriptomics based point-clouds (Haviv et al., 2024b, ).
During embryogenesis, specific regions within the primitive gut tube differentiate into organs such as the liver or lungs based on interactions between the gut and surrounding mesenchyme (Nowotschin et al.,, 2019). Applied on environments of gut-tube cells from a seqFISH dataset of mouse embryogenesis (Lohoff et al.,, 2022), WFM synthesized cellular niches conditioned on organ labels, thus demanding an understanding of the interplay between spatial context and phenotype. Despite the intricate nature of the gastrulation process, compunded by the dataset’s dimensionality, WFM can accurately generate organ-specific niches; see Figure 6 and Table 2.
Figure 6: The niche of each cell is the point-cloud from gene-expression profiles for cells in its environments. Using the top 16 principal components of gene expression, WFM learns to generate high-dimensional cellular microenvironments of gut-tube cells conditioned on the gastrulating organs.
@@ 5 Conclusion and outlook
This work shows how to appropriately lift the Riemannian flow matching paradigm of Chen and Lipman, (2023) to the Wasserstein space, resulting in Wasserstein flow matching. Our motivations stem from modern datasets, where each sample of data can itself be viewed as a probability distribution, necessitating this extension for generative modeling purposes. Our contributions are algorithmic in nature, which incorporate various elements, such as estimating optimal transport maps via entropic optimal transport, closed-form expressions over the Bures–Wasserstein space, and attention mechanisms in neural network architectures. Our algorithm is capable of generating realistic data from Gaussian and variable-size or high-dimensional point-clouds. Both contexts are highly relevant in single-cell and spatial transcriptomics for synthesizing of microenvironments and cellular states.
Acknowledgements
@@ Appendix A Entropic estimation of OT maps
We briefly discuss how to estimate optimal transport maps between point-clouds using entropic optimal transport. We refer the interested reader to Pooladian and Niles-Weed, (2021) for more information on this approach.
We first outline the numerical aspects of the approach; we follow Peyré and Cuturi, (2019). Let $\mu=\sum_{i}m^{-1}\delta_{x_{i}}$ and $\nu=\sum_{j}n^{-1}\delta_{y_{j}}$, where $\bm{X}=\{x_{1},\ldots,x_{m}\}$, $\bm{Y}=\{y_{1},\ldots,y_{n}\}$. We first define the following polyhedral constraint set
@@
where $\varepsilon>0$, $H(P)\coloneqq\sum_{i,j}P_{i,j}(\log(P_{i,j})-1)$, and $C_{i,j}\coloneqq\|x_{i}-y_{j}\|^{2}_{2}$. Sinkhorn’s matrix scaling algorithm (Sinkhorn,, 1964) makes it possible to solve for $\bm{P}^{\star}$ with a runtime of $O(mn/\varepsilon)$ (Altschuler et al.,, 2017). We briefly stress two points:
@@
    When $\varepsilon=0$, the objective becomes a standard linear program with a runtime of $\tilde{O}(mn(m+n))$ (up to log factors) (Peyré and Cuturi,, 2019, Chapter 3). However, this approach does not come with a GPU-efficient implementation and we decide to omit it entirely.
In all our experiments, we used the open-source package OTT-JAX [^2] See https://ott-jax.readthedocs.io/en/latest/. to compute the entropic coupling and the out-of-sample mapping (Cuturi et al.,, 2022).
@@ A.2 Entropic transport map: An out-of-sample estimator
A primal-dual relationship of the strictly convex program equation 13 shows that there exist vectors $(\bm{f}^{\star},\bm{g}^{\star})\in\mathbb{R}^{m}\times\mathbb{R}^{n}$ such that
@@
These two vectors are called the Kantorovich potentials, which are initially defined on the support of $\mu$ and $\nu$, respectively. However, they can be readily extended to all of $\mathbb{R}^{d}$ (Mena and Niles-Weed,, 2019), resulting in two functions
@@
Following Pooladian and Niles-Weed, (2021), we can define the entropic transport map, where the last equality is a simple calculation:

### v2

@@ 4.1 Families of Gaussians
seqFISH (16D) $\mathbf{3.11\cdot 10^{-1}}$ $5.31\cdot 10^{-1}$ $6.82\cdot 10^{-1}$
@@
Table 2: Average min. $W^{2}_{2}$ distance between each generated Gaussian and the reference datasets. Despite identical training schemes, BW-FM (R) outperforms other approaches on both synthetic and real data across several dimensions.
In all cases, we assess the quality of the learned flows by computing the the minimum distance between each generated Gaussian and the dataset using the (squared) $2$-Wasserstein distance; see Table 2. Notably, the flows generated by Frobenius FM and BW-FM (Euclidean) do not strictly adhere to the geometry of the Bures-Wasserstein manifold, requiring synthesized covariance matrices to be projected onto the space of PSD matrices. In contrast, the Riemannian BW-FM algorithm consistently produces valid and accurate results across all dimensions and datasets.
4.1.1 Toy Datasets
As a first test, we design a dataset of Gaussians centered on a spiral; See Figures 2 & S1. When there were only few samples, only Riemannian BW-FM reconstructed the data, despite other benchmark methods following identical training regimes. On the complete 128-sample dataset, generalizes beyond the training data and synthesizes novel Gaussians lying on the spiral. BW-FM shares this generalization feature with standard FM, and is able to learn the structure underlying the measure from the training data.
@@ 4.1.2 Single-Cell Genomics
Spatial transcriptomics are a set of techniques which build on single-cell genomics and preserve physical information of cells’ location in tissues and assay their gene expression. Haviv et al. (2024b) demonstrated that a cell’s microenvironment can be effectively characterized using the mean and covariance of the surrounding cells gene expression. This representation captures key features of cellular neighborhoods and transforms spatial transcriptomics datasets into a measure within the Bures–Wasserstein space, highlighting the value of generative modeling in this context.
During embryogenesis, specific regions within the primitive gut tube differentiate into organs such as the liver or lungs based on interactions between the gut and surrounding mesenchyme (Nowotschin et al., 2019). Applied on environments of gut-tube cells from a seqFISH dataset of mouse embryogenesis (Lohoff et al., 2022), BW-FM synthesized Gaussian cellular niches conditioned on organ labels, thus demanding an understanding of the interplay between spatial context and phenotype. Despite the intricate nature of the gastrulation process, compounded by the dataset’s dimensionality, BW-FM can accurately generate organ-specific niches; see Figure S2.
Figure 4: Generated distributions from MNIST and Letters datasets using WFM. Each distribution is realized as a point-cloud with naturally varying sample sizes ($n$), demonstrating WFM’s unique ability to learn from and generate distributions without requiring fixed-size realizations.
@@
Another common instance of BW manifolds arising in single-cell genomics is through aggregating cells into common states. These clusters can be summarized by their mean gene expression and its covariance. On a scRNA-seq atlas elucidating human immune response to COVID (Stephenson et al., 2021), we combine cells into MetaCells (Persad et al., 2023), and quantify the gene expression mean and covariance for each. Here too we apply BW-FM conditioned on cell-state, which encompass the heterogeneity of immune profiles appearing as response to COVID infection. Despite the plurality of labels, BW-FM can synthesize correct examples for each condition; see Figure S3.
4.2 Families of general distributions
Recall that for general distributions, we use entropic optimal transport to estimate McCann interpolations and velocities (Cuturi, 2013). With high-resolution CAD datasets, we samplr 2048-point realizations from each shape to approximate OT maps during training. For spatial transcriptomics, MNIST and Letters datasets, we work directly with their naturally varying sample sizes, leveraging entropic OT’s ability to handle unequal point-cloud sizes while maintaining accurate transport estimation. These align with our assumption that the underlying distributions are continuous, with point-clouds serving as discrete realizations.
@@ 4.2 Families of general distributions
WFM (ours) 69.88 64.44 57.62 57.93 53.41 58.10
Table 3: Performance evaluation using 1-NN Accuracy (Earth Mover’s Distance and Chamfer Distance) on distributions from ShapeNet, realized as uniform 3D point-clouds. Top three results for each metric are highlighted (1st, 2nd, 3rd). WFM achieves state-of-the-art CD performance on 2 datasets while maintaining competitive results across other metrics (data from Wu et al. (2023a)), while producing diverse samples, see Figure 3.
We compare WFM to many other generative models. Following in their footsteps, we measure generation quality based on the 1-Nearest-Neighbour accuracy metric between generated and test-set shapes. On uniform, 3D datasets, WFM is competitive with current approaches, but exemplifies itself with its unique ability to generate samples with varying sizes and in high-dimensions; see Tables 3 & 4.
4.2.1 2D & 3D Shapes
Derived from 3D CAD designs, ShapeNet & ModelNet (Wu et al., 2015; Chang et al., 2015) are touchstone shape datasets in computational geometry. Trained individually on samples from the chair, car and plane classes of ShapeNet, WFM synthesized high quality shapes with diverse profiles and matches the performance of previous 3D generation algorithms; see Figure 3 and Table 3. Our framework’s versatility allows for seamless integration of label information during training, enabling the synthesis of shapes conditioned on specific classes. On the full 40-class ModelNet dataset, WFM learned condition-dependent flows between distributions, enabling generation of diverse shapes from the same initial distribution based on the desired label; see Figure 3. We stress that WFM is not restricted to only noisy source measures but can generate transformations between any two collections of distributions. To this end, we demonstrate that WFM can interpolate between two arbitrary elements in the dataset (e.g., between a lamp and a handbag) and complete shapes based on partial profiles (e.g., generate the remaining parts of a plane); see Figure S5.
Current methods WFM (ours)
CD $\downarrow$ EMD $\downarrow$ CD $\downarrow$ EMD $\downarrow$
Variable-sized realizations
MNIST ($4$) NA NA 63.34 59.97
Letters ($A$) NA NA 62.12 58.68
High-dimensional distributions
MERFISH NA NA 52.80 54.52
XENIUM NA NA 60.69 64.20
Table 4: 1-Nearest-Neighbour Accuracy demonstrating WFM’s unique capabilities. Using transformers and entropic transport maps, WFM handles both variable-sized realizations and high-dimensional distributions, unlike previous approaches limited to fixed-size, low-dimensional samples.
Another novel facet of WFM is its ability to perform generative modeling from inhomogeneous datasets, where the number of points per point-cloud realization varies between independent samples. This happens in the MNIST or Letters datasets, where data is generated by thresholding low-resolution grayscale images. WFM sets itself apart from other methods, which are restricted to uniform datasets, by leveraging the entropic OT map’s ability to compute feasible transformations between empirical samples of different sizes. Our experiments in Figure 4 demonstrate that WFM generates high-quality & diverse samples, despite large variability in the number of particles, itself a novel contribution.
4.2.2 Spatial Transcriptomics
@@
In spatial transcriptomics, a cell’s niche is represented by the distribution of gene expression in its neighboring cells, which we observe through its immediate nearest neighbors in high-dimensional space. This approach is complementary to the BW representation of a niche (recall Section 4.1.2), and serves as a more high fidelity view suited for fine-grain interactions. Due to their high dimensionality, these cellular microenvironment distributions have remained beyond the reach of generative models that depend on voxel-based neural networks. Instead, WFM uses transformers, whose permutation equivariance and indifference to dimensionality make them natural architectures for modeling these high-dimensional distributions through their point-cloud realizations (Haviv et al., 2024b).
Figure 5: WFM enables high-dimensional generation of cellular microenvironments. (A) Wasserstein 2D embedding of observed Sst-neuron niches in motor cortex, colored by subtype. (B) WFM generated niches faithfuly reproduce of the tissue landscape
In the motor cortex, somatostatin (Sst) interneurons are organized into phenotypically distinct subtypes, each localized to specific cortical layers (Wu et al., 2023b). Using the 254-gene MERFISH atlas (Zhang et al., 2021), we represent each Sst neuron’s microenvironment as a distribution in gene expression space, observed through its 16 nearest neighbors. Despite operating in this high-dimensional space, WFM successfully generates niche distributions that capture both global tissue organization and subtype-specific molecular signatures; see Figures 5 & 6.
Figure 6: Examples of generated niches for Calb2+ and Th+ Sst-neurons (see Figure 5) display expected gene-expression signatures, with Calb2+ niches enriched for upper-layer markers (Cux2) and Th+ for deep-layer markers (Ntng2 and Foxp2). Each niche is represented as a high-dimensional distribution realized through point-clouds, demonstrating WFM’s unique capability to generate distributions in spaces where previous methods, limited by voxelization, cannot operate.
@@ 5 Conclusion and outlook
This work shows how to appropriately lift the Riemannian flow matching paradigm of Chen & Lipman (2023) to the Wasserstein space, resulting in Wasserstein flow matching. Our motivations stem from modern datasets, where each data sample is a probability distribution, necessitating this extension for generative modeling purposes. Our contributions are algorithmic in nature, which incorporate various elements, such as estimating optimal transport maps via entropic OT, closed-form expressions over the Bures–Wasserstein space, and attention mechanisms in neural network architectures. Our algorithm is capable of generating realistic data from Gaussian and general distribution realized as variable-size or high-dimensional point-clouds. Both contexts are highly relevant in single-cell transcriptomics for synthesizing of microenvironments and cellular states.
6 Impact Statement
Wasserstein Flow Matching is designed for generative modeling of distributions, with applications in computational biology. As with any generative model, care should be taken when using WFM to draw biological conclusions.
7 Acknowledgements
@@ Appendix A Entropic estimation of OT maps
We briefly discuss how to estimate optimal transport maps between distributions using entropic OT. We refer the interested reader to Pooladian & Niles-Weed (2021) for more information on this approach.
We first outline the numerical aspects of the approach; we follow Peyré & Cuturi (2019). Let $\mu=\sum_{i}m^{-1}\delta_{x_{i}}$ and $\nu=\sum_{j}n^{-1}\delta_{y_{j}}$ be two empirical distributions where $\bm{X}=\{x_{1},\ldots,x_{m}\}$, $\bm{Y}=\{y_{1},\ldots,y_{n}\}$. We first define the following polyhedral constraint set
@@
where $\varepsilon>0$, $H(P)\coloneqq\sum_{i,j}P_{i,j}(\log(P_{i,j})-1)$, and $C_{i,j}\coloneqq\|x_{i}-y_{j}\|^{2}_{2}$. Sinkhorn’s matrix scaling algorithm (Sinkhorn, 1964) makes it possible to solve for $\bm{P}^{\star}$ with a runtime of $O(mn/\varepsilon)$ (Altschuler et al., 2017). We briefly stress three points:
@@
    When $\varepsilon=0$, the objective becomes a standard linear program with a runtime of $\tilde{O}(mn(m+n))$ (up to log factors) (Peyré & Cuturi, 2019, Chapter 3). While we include a CPU implementation (Flamary et al., 2021) in the WFM codebase, this approach lacks GPU efficiency and substantially increases training time, making it impractical for most use cases.
  3. 3.
    Instead, the regularization parameter $\varepsilon$ serves as a tunable training hyperparameter. Lower $\varepsilon$ values better approximate true the optimal transport map but require more Sinkhorn iterations for convergence, creating a direct trade-off between accuracy and computational efficiency.
In all our experiments, we used the open-source package OTT-JAX (See https://ott-jax.readthedocs.io/en/latest/) to compute the entropic coupling and the out-of-sample mapping (Cuturi et al., 2022).
@@ A.2 Entropic transport map: An out-of-sample estimator
A primal-dual relationship of the strictly convex program (12) shows that there exist vectors $(\bm{f}^{\star},\bm{g}^{\star})\in\mathbb{R}^{m}\times\mathbb{R}^{n}$ such that
@@
These two vectors are called the Kantorovich potentials, which are initially defined on the support of $\mu$ and $\nu$, respectively. However, they can be readily extended to all of $\mathbb{R}^{d}$ (Mena & Niles-Weed, 2019), resulting in two functions
@@
Following Pooladian & Niles-Weed (2021), we can define the entropic transport map, where the last equality is a simple calculation: